A patient transport vehicle is running at a speed of 80 m/s. It is the decelerated at the rate of 10m/s. When will it stop?

1
Expert's answer
2021-12-27T08:01:16-0500

acceleration:

"a=\frac{v_2-v_1}{t}"


we have:

"a=-10" m/s2

"v_2=0"

"v_1=80" m/s


vehicle will stop after:

"t=\frac{v_2-v_1}{a}=\frac{80}{10}=8" s